When Tim started football 3 years ago, his coach was twice his age. Currently, Tim's coach is 10 years older than him. How old is Tim now?

Answer by Maths68(1474)

Let
Present age of Time = t
Present age of Coach= c
Three years ago means we have to subtract 3 from their current ages
c-3=t-3
Then
Coach was twice the age to time
c-3=2(t-3)
c-3=2t-6
c=2t-6+3
c=2t-3...........(1)
Tim's coach is 10 years older than him means if we add 10 to Tim's age Coach and Time will be of equal age
c=t+10.........(2)
Put the value of c from (2) to (1)
c=2t-3...........(1)
t+10=2t-3
t+10=2t-3
t=2t-3-10
t-2t=-13
-t=-13
Multiply by (-1) both sides of the equation
(-1)(-t)=(-1)(-13)
t=13
Put the value of t in (2)
c=t+10.........(2)
c=13+10
c=23
Present age of Time = t = 13
Present age of Coach= c = 23